A group of students is determining the density of a cube of metal. Using a calipers, they measure all three sides of the cube to be 2.000+ .002 cm . What is the volume they calculate, including the uncertainty?
Answer #1

The length of the side of the cube is

a = (2.000 + 0.002) cm

The volume of the cube is

V = n = 2.000* = 8 cm

And the uncertainty in the calculation of the volume can be given by

u(V) 3u(a)

31 = u(V) = u(a) = -Xºx 0.002 = 0.024 cm

So, the volume of the cube with the uncertainty is

V = (8.000 + 0.024) cm3